# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

### v2

@@ 5 RAE-2822 airfoil computer experiment
Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9]$
We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$. Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Once we halt contour location (Section 3.1), we use entropy to inform our hybrid MC estimator (Section 3.2). Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we opt to repeat this entire exercise twice, with a newly randomized starting design. We consider the agreement between the two randomized repetitions as an indication that our procedure is well-calibrated.
@@
Figure 7 reports the progress in $\hat{\alpha}$ over the course of each design. Our stopping criterion halted the sequential designs at $N=180$ and $N=160$, leaving $B=320$ and $B=340$ for the second design stage, respectively. Our two hybrid Monte Carlo estimates were $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$. These values are indicated by the red dashed lines in Figure 7. The grey shading shows $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$, calculated from these final estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
@@ 6 Discussion
We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments. Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator. Then we use the remaining budget of simulator evaluations on the Monte Carlo samples with highest entropy, to ultimately return a hybrid MC estimate of $\alpha$. Our design strategically leverages the uncertainty quantification provided by the surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. While we limited our experiments to GP and DGP surrogates, our proposed scheme is applicable to any surrogate which provides posterior predictive distributions (which are required for entropy calculations). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. Posterior predictions must be computed at all $M$ samples multiple times. While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial. Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason. We could tackle smaller probabilities, but it would require larger $M$, which would eventually become impractical. Although we are focused on expensive simulators whose computational costs overshadow the cost of these surrogate predictions, our approach may simply not be suited to extremely rare failure probabilities.
@@
In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted. If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met. While we prefer to have additional evaluations reserved for the estimation stage, we contend that using the entire budget for contour location to provide a surrogate MC estimate is the best option in this scenario. As we’ve shown, importance sampling based methods perform poorly when data is limited.
@@
We have exclusively focused on deterministic computer experiments. In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design. Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$. Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate. Different design choices could be warranted for stochastic simulators (Baker et al.,, 2022). In particular, the hybrid estimator of Eq. (7) should not make use of noisy $f(\mathbf{x})$. Replicates in the design may be crucial (Binois et al.,, 2019).
@@
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will never occur. Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as the black open circles around the edges of Figure 3). We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored.
@@
Estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization. For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold. The ability to effectively estimate failure probabilites from limited data, which we have tackled here, is integral to the success of this larger enterprise.
@@
The Plateau function (Booth et al.,, 2024b) is defined in arbitrary dimension over $[0,1]^{d}$ as

### v3

@@ 5 RAE-2822 airfoil computer experiment
Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9].$
We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$. Booth et al., (2024c) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location (Stage 1) following Booth et al., (2024c), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is within the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Once we halt contour location (Section 3.1), we proceed to Stage 2, using entropy to inform our hybrid MC estimator (Section 3.2). Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we opt to repeat this entire exercise twice, with a newly randomized starting design. We consider the agreement between the two randomized repetitions as an indication that our procedure is well-calibrated.
@@
Figure 7 reports the progress in $\hat{\alpha}$ over the course of each design. Our stopping criterion halted the sequential designs at $n=180$ and $n=160$, leaving $B=320$ and $B=340$ for the second design stage, respectively. Our two hybrid Monte Carlo estimates were $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$. These values are indicated by the red dashed lines in Figure 7. The grey shading shows $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$, calculated from these final estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator of Stage 2 is effectively adjusting for surrogate inaccuracies remaining after Stage 1.
@@ 6 Discussion
We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments, which outperforms both exhaustive contour location and surrogate-informed importance sampling. Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator. Then we use the remaining budget of simulator evaluations on the Monte Carlo samples with highest entropy, to ultimately return a hybrid MC estimate of $\alpha$. Our design strategically leverages the uncertainty quantification provided by the surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. While we limited our experiments to GP and DGP surrogates, our proposed scheme is applicable to any surrogate which provides posterior predictive distributions (which are required for entropy calculations). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial. The computational cost of obtaining these predictions increases with $n/N$, $d$, and $M$, but it also depends on the surrogate and computing environment. For context, the most expensive predictions we needed in our exercises were those of our motivating airfoil simulator, with the fully-Bayesian DGP trained on all $N=500$ points in 7-dimensions. With parallelization over 10 CPUs, it took just over 10 minutes to obtain one million predictions. For smaller $\alpha$ requiring $M$ in the billions, the cost of these surrogate predictions may present an obstacle, but it is still far cheaper than comparative evaluations of the true expensive computer simulator.
There may be avenues to circumvent surrogate predictions at all $X_{M}$ locations. We hypothesize that a classification rule based on nearest neighbors could work well: for a particular $\mathbf{x}_{i}$, if the surrogate predicts that all of its $k$-many nearest neighbors are “passes,” then we could presume $\mathbbm{1}_{\{\mu_{n}(\mathbf{x}_{i})>t\}}=0$ without needing to obtain $\mu_{n}(\mathbf{x}_{i})$ directly. This requires effective tuning of $k$, and would only speed-up computations if the cost of obtaining the nearest neighbor sets is less than that of obtaining the single surrogate prediction. For these reasons, we did not entertain such workarounds in our current work.
@@
In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted. If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met. While we prefer to have additional evaluations reserved for Stage 2, we contend that using the entire budget for contour location in Stage 1 to inform a traditional surrogate MC estimate is the best option in this scenario. As we’ve shown, importance sampling based methods perform poorly when data is limited.
@@
We have exclusively focused on deterministic computer experiments. In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design. Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations. Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate. Stochastic simulators (e.g., Baker et al.,, 2022) present a unique challenge as repeated observations of noisy $f(\mathbf{x})$ may result in a mix of passes and failures for the same $\mathbf{x}$. To accommodate this, we could have broadened the formulation of Eq. (1) by replacing the indicator function $\mathbbm{1}_{\{f(\mathbf{x})>t\}}$ with the unknown probability $\mathrm{P}(f(\mathbf{x})>t)$, which recovers the indicator when $f$ is deterministic. The key advantage of our hybrid estimator – that observing $f(\mathbf{x}_{i})$ ensures correct classification of $\mathbbm{1}_{\{f(\mathbf{x}_{i})>t\}}$ – is not applicable to noisy simulators. Instead, new active learning strategies that strategically allocate replicates (Binois et al.,, 2019) and inform a MC estimator leveraging surrogate posterior predictive probabilities are warranted. We leave this as an avenue for future work.
@@
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will never occur. Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as those around the edges in Figure 2). We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored. Furthermore, while we focused on stochastic inputs, it is possible to extend our approach to incorporate controllable/design variables by using point masses in $p(\mathbf{x})$.
@@
Finally, estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization. For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold. The ability to effectively estimate failure probabilities from limited data, which we have tackled here, is integral to the success of this larger enterprise. While we focused on allocating a predetermined budget of evaluations, it could be useful to develop a decision rule for halting Stage 2, potentially reserving some evaluations, particularly if this design is within a larger optimization loop.
@@ Appendix A Test functions
The Plateau function (Booth et al.,, 2024c) is defined in arbitrary dimension over $[0,1]^{d}$ as
@@
Appendix B Batch size comparison
Here we consider variations of the batch sizes used within the second stage of our design. When Stage 1 is complete and $n$-many evaluations have been spent on contour location, the remaining budget ($B=N-n$) is to be spent on the highest entropy samples from $X_{M}$. Since the surrogate is used to select these samples, there is an option to select them in smaller batches, updating the surrogate with the new observations between each batch (indicated by the dotted arrow in Figure 1). But smaller batches require more computation as surrogate predictions must be obtained at all $X_{M}$ locations after every batch update.
To investigate, we consider three different batch sizes for the Herbie function example of Section 4. First, we choose all high entropy points in a single batch of size $B$ (this is the method used throughout the paper, thus these results are duplicated from Figure 6). Second, we consider selecting one high entropy point at a time, updating the surrogate after every high entropy point is observed to better inform the selection of subsequent high entropy points. Third, we consider selecting high entropy points in batches of size 10. Note, changing the batch size used within Stage 2 has no effect on Stage 1; each of these batch variations starts in the same place after the completion of contour location and uses the same $X_{M}$.
Figure 8 shows the resulting failure probability estimates across 30 Monte Carlo repetitions. There is no significant difference in performance among the three batch sizes. The grey shaded region marks $\{\alpha\pm 2\sigma_{\alpha}\}$; it indicates the level of variability that we would expect from any Monte Carlo estimator of specified $M$, even if every sample was classified accurately. All methods consistently fall within this region, indicating effective estimation (note the narrower y-axis range compared to Figure 6). Given the computational expense of repeatedly updating the surrogate, we embrace the single batch approach.
Figure 8: Estimated failure probabilities shown over the true failure probability (dashed red) ±$\pm 2$ standard errors (grey shading) for proposed two-stage designs of the Herbie function. Variations use different batch sizes for selecting the highest entropy samples within Stage 2.